### v1

# Inertial antiferromagnetic resonance driven by spin-orbit torques

Abstract
It is widely accepted that the handedness of a resonant mode is an intrinsic property. We show that, by tailoring the polarization and handedness of alternating spin-orbit torques used as the driving force, the polarization state and handedness of inertial resonant modes in an antiferromagnet can be actively controlled. In contrast to ferromagnets, whose resonant-mode polarization is essentially fixed, antiferromagnetic inertial modes can continuously evolve from elliptic through circular to linear polarization as the driving polarization is varied. We further identify an inertia-dependent critical degree of driving polarization at which the mode becomes linearly polarized while its handedness reverses.
I introduction
Spin inertia [1] in magnetically ordered materials can be induced by various microscopic mechanisms, as demonstrated by recent studies. First-principles spin dynamics have shown that both damping and inertia emerge consistently from the electronic structure, emphasizing the role of spin-orbit coupling in dissipative and inertial spin dynamics [2, 3]. A complementary approach based on spin-bath coupling reveals that a non-Markovian memory kernel yields an effective inertial term in the spin equation of motion [4, 5]. Additionally, nonadiabatic interactions with environmental degrees of freedom can produce inertia-like corrections [6]. A classical interpretation models the spin as a circular current loop, leading to inertial terms from the underlying mechanical dynamics [7, 8].
Spin inertia gives rise to magnetic nutation [9, 10, 6, 11, 12], which manifests as nutation resonance [13, 14, 15, 16]. Furthermore, inertia modifies the self-oscillatory dynamics of magnetization in both antiferromagnets (AFMs) [17, 18] and ferromagnets (FMs) [19]. The characteristic influence of spin inertia on these uniform magnetic dynamics can serve as an additional experimental probe for its detection.
Apart from experimental observations of nutation [13, 14, 15, 16], extensive efforts have been invested in the theoretic research of nutational resonance [20, 21, 22, 26, 27, 29, 28, 23, 24, 25]. It was found that the inertial magnetization dynamics are described by the inertial Landau-Lifshitz-Gilbert (ILLG) equation, which predicts a nutation resonance in addition to conventional ferromagnetic resonance. Olive et al. [20, 21] numerically demonstrated a secondary high-frequency peak whose amplitude and position depend on the inertial relaxation time, the damping, and the applied field. Subsequent analytical studies quantified shifts of resonant frequencies and linewidths of both precession and nutation, revealing a non-trivial dependence on damping and magnetic anisotropy [22, 24, 23, 25]. Another analytical study predicts that these effects are enhanced in AFMs due to exchange interaction [26]. Remarkably, when driven by a circularly polarized magnetic field, FMs exhibit a single nutational peak and a single precessional peak, whereas AFMs and ferrimagnets present these peaks in pairs. In the latter two systems, inter- and intra-sublattice couplings further tune these resonances, indicating distinct mechanisms in multi-sublattice systems [27, 28]. Finally, spin pumping at terahertz nutation frequencies produces a spin current opposite in direction to that generated by precessional modes, providing an unambiguous experimental signature of nutation [29].
We note that most studies on inertial resonances have primarily focused the resonant frequency, determined by the intrinsic magnetic properties. However, the polarization of the resonant mode, which depends on both the magnetic material and the external drive, remains comparatively unexplored, despite offering an extra degree of freedom – in addition to amplitude and phase – for information encoding [30]. Experimentally, resonance is usually driven by circularly or linearly polarized magnetic fields. Accordingly, the nutational resonances in the current theoretical studies were assumed to be excited by circularly [22, 26, 27, 29, 28, 25] or linearly polarized [20, 21] oscillating fields.
An alternative utilises current-induced spin-torque resonance, realised via spin-transfer torques in spin-valve structures [31, 32, 33, 34, 35, 36] and spin-orbit torques (SOTs) in heavy-metal/FM bilayers [37, 38, 39, 40], both of which are efficient and scalable in spintronic devices. In both schemes an alternating charge current generates an oscillating spin current that exerts a spin torque on the magnetization. Whereas spin-transfer torque employs a current perpendicular to the film, spin-orbit torque relies on an in-plane current, enabling finer control of the polarization (linear, elliptical or circular) and handedness (left- or right-handed) of the driving torque and, consequently, of the resonant mode.
Motivated by the above considerations, we present the investigation of inertial resonant modes in AFMs driven by alternating spin-orbit torques, with an emphasis on their polarization and handedness.
II model
As illustrated in Fig. 1, we study an inertial AFM film attached to a heavy metal (HM) with strong spin-orbit coupling, driven by oscillating in-plane currents. The bipartite AFM dynamics, that includes magnetic inertia, Gilbert damping, and SOTs, are captured phenomenologically by two exchange-coupled ILLG equations. For sublattice $k$, the ILLG equation reads:
$\frac{d\mathbf{m}_{k}}{dt}=\mathbf{m}_{k}\times\frac{d\mathcal{E}}{d\mathbf{m}_{k}}+\alpha\mathbf{m}_{k}\times\frac{d\mathbf{m}_{k}}{dt}+\eta\mathbf{m}_{k}\times\frac{d^{2}\mathbf{m}_{k}}{dt^{2}}+\boldsymbol{\tau}_{k},$ (1)
where $\mathbf{m}_{k}$ is the unit magnetization vector of sublattice $k$ ($k=1,2$), $\alpha$ is the Gilbert damping constant, and $\eta$ is the inertial relaxation time, typically ranging from femtoseconds to picoseconds, as predicted by ab initio calculations [41] and confirmed experimentally [13, 14, 15]. The magnetic energy, arising from the magnetocrystalline anisotropy and the intersublattice exchange interaction, is expressed in frequency units as
$\mathcal{E}=\omega_{E}\mathbf{m}_{1}\cdot\mathbf{m}_{2}+\omega_{K}\sum_{k}\left(\mathbf{m}_{k}\cdot\mathbf{e}_{z}\right)^{2},$ (2)
where $\omega_{E(K)}=\gamma_{0}H_{E(K)}$, with $H_{E}$ and $H_{K}$ being the exchange field and the anisotropy field, respectively. $\gamma_{0}=g\mu_{0}\mu_{B}/\hbar$ is the gyromagnetic ratio with $g$ being the Landé $g$-factor, $\mu_{0}$ the vacuum susceptibility, $\mu_{B}$ the Bohr magneton, and $\hbar$ the reduced Plank constant. The uniaxial anisotropy axis is oriented perpendicular to the film plane.
The last term in Eq. (1) represents the SOTs, originating from spin transfer between the local magnetization and spin currents generated in the adjacent HM layer. Including both dampinglike and fieldlike components [42], the SOTs are given by
$\mathbf{\tau}_{k}=-\rho\left\{\mathbf{m}_{k}\!\times\!\left[\mathbf{m}_{k}\!\times\!\left(\mathbf{e}_{z}\!\times\!\mathbf{j}_{e}\right)\right]\!+\!\beta\mathbf{m}_{k}\!\times\!\left(\mathbf{e}_{z}\!\times\!\mathbf{j}_{e}\right)\right\},$ (3)
where $\beta$ is the ratio of fieldlike to dampinglike SOT. The prefactor $\rho=\xi\mu_{B}/(eM_{s}d)$ characterizes the SOT strength per unit current density, with $M_{s}$ being the saturation magnetization of sublattice, $d$ the thickness of the AFM layer, and $e$ the elementary charge. The SOT efficiency is defined as $\xi=T_{\mathrm{int}}\theta_{\mathrm{sh}}$ [43, 44], where $\theta_{\mathrm{sh}}$ is the spin Hall angle [42] and $T_{\mathrm{int}}$ the spin transparency of the HM/AFM interface [45]. The in-plane electric current density $\mathbf{j}_{e}$ is expressed in components as
$\mathbf{j}_{e}=j_{x}\mathbf{e}_{x}+j_{y}\mathbf{e}_{y}.$ (4)
To excite stationary modes, in-plane alternating electric currents are applied, as shown in Fig. 1. For analytical convenience, the current is expressed in complex form $(j_{x},j_{y})=(J_{x},iJ_{y})e^{-i\omega t}$, where $J_{x(y)}>0$ denote the current amplitudes. In this representation, the phase of $j_{y}$ leads (lags) that of $j_{x}$ by $90^{\circ}$ for positive (negative) $\omega$, corresponding to right-handed (left-handed) polarized SOTs. Taking the real part yields the physical current: $J_{x}\cos\omega t\,\mathbf{e}_{x}+J_{y}\sin\omega t\,\mathbf{e}_{y}$.
Figure 1: (color online). Schematic diagram of model.
III resonant spectrum
Due to the uniaxial anisotropy and intersublattice exchange [Eq. (2)], the static equilibrium magnetizations of the two sublattices align along the ±$\pm z$ directions, respectively. For small deviations from equilibrium, the $z$-components of the magnetizations remain unchanged to first order, such that $m_{1}^{z}\approx 1$ and $m_{2}^{z}\approx-1$. To describe the magnetic response of the AFM in the linear regime, we adopt the ansatz $\mathbf{m}_{k}=(-1)^{k+1}\mathbf{e}_{z}+m_{k}^{x}(t)\mathbf{e}_{x}+m_{k}^{y}(t)\mathbf{e}_{y}$, where $m_{k}^{x}(t)$, $m_{k}^{y}(t)$ are complex-valued functions encoding the amplitude and phase of the response to SOTs. Inserting the ansatz into Eq. (1), assuming a harmonic time dependence of $m_{k}^{x}(t)$, $m_{k}^{y}(t)$ (i.e., replacing $d/dt$ by $-i\omega$), and keeping only terms linear in $m_{k}^{x}(t)$, $m_{k}^{y}(t)$, $j_{x}$ and $j_{y}$, one obtains a four-dimensional linear inhomogeneous system of ordinary differential equations that describes damped magnetic oscillations driven by periodic forces. Solving this system yields the steady-state linear response modes:
$\left(\begin{array}[]{c}m_{1}^{x}(t)\\ m_{1}^{y}(t)\\ m_{2}^{x}(t)\\ m_{2}^{y}(t)\end{array}\right)=\chi\left(\begin{array}[]{c}j_{x}(t)\\ j_{y}(t)\\ j_{x}(t)\\ j_{y}(t)\end{array}\right).$ (5)
Here, in analogy with the Polder susceptibility tensor [46] describing the linear response to an alternating magnetic field, we define a susceptibility tensor characterizing the linear response to alternating SOTs, which reads
$\chi=\frac{\rho}{\Delta}\left(\begin{array}[]{cccc}\chi_{1}&\chi_{2}&-\omega_{E}&\beta\omega_{E}\\ -\chi_{2}&\chi_{1}&-\beta\omega_{E}&-\omega_{E}\\ \omega_{E}&\beta\omega_{E}&-\chi_{1}&\chi_{2}\\ -\beta\omega_{E}&\omega_{E}&-\chi_{2}&-\chi_{1}\end{array}\right),$ (6)
where $\chi_{1}=(\eta\omega^{2}-\omega_{E}-\omega_{K})+i(\alpha-\beta)\omega$, $\chi_{2}=\beta(\eta\omega^{2}-\omega_{E}-\omega_{K})+i(1+\alpha\beta)\omega$, and
$\Delta=\left(\Omega_{1}-i\alpha\omega\right)\left(\Omega_{2}-i\alpha\omega\right)-\omega^{2},$ (7)
with
$\Omega_{1}$ $=$ $2\omega_{E}+\omega_{K}-\eta\omega^{2},$ (8)
$\Omega_{2}$ $=$ $\omega_{K}-\eta\omega^{2}.$ (9)
To analyze the properties of the steady-state modes, it is convenient to take the real part of Eq. (5). The resulting harmonic oscillations of the magnetization driven by alternating SOTs with frequency $\omega$ are given by
$m_{k}^{x(y)}=\mathcal{A}_{k}^{x(y)}\cos\left(\omega t+\phi_{k}^{x(y)}\right).$ (10)
The polarization degree of the linear modes is set by the ratio $\mathcal{A}_{k}^{y}/\mathcal{A}_{k}^{x}$, while the phase difference between $m_{k}^{x}$ and $m_{k}^{y}$ determines the handedness of $\mathbf{m}_{k}$. In Eq. (10) the amplitudes are
$\mathcal{A}^{x(y)}_{1,2}=\rho\sqrt{\frac{P_{1}J^{2}_{x(y)}+P_{2}J^{2}_{y(x)}\pm P_{3}J_{x}J_{y}}{\widetilde{\Delta}}},$ (11)
where upper and lower signs represent sublattices $1$ and $2$, respectively. This sign convention is maintained throughout unless stated otherwise. To simplify the expressions, we introduce $P_{1}=(\alpha-\beta)^{2}\omega^{2}+\Omega_{1}^{2}$, $P_{2}=(1+\alpha\beta)^{2}\omega^{2}+\beta^{2}\Omega_{2}^{2}$, $P_{3}=2\omega[2(1+\alpha\beta)\omega_{E}+(1+\beta^{2})\Omega_{2}]$, and
$\widetilde{\Delta}=\left[\left(1\!-\!\alpha^{2}\right)\omega^{2}\!-\!\Omega_{1}\Omega_{2}\right]^{2}\!+\!4\alpha^{2}\omega^{2}\left(\omega^{2}\!+\!\omega_{E}^{2}\right).$ (12)
Figure 2: (color online). Resonant spectra for various inertia parameters $\eta$. Peaks enclosed in the circles (ellipses) correspond to precessional (nutational) resonances. Both axes are shown on a logarithmic scale; low-frequency parts are omitted for clarity. Curves are evaluated from Eq. (11) using $\omega_{E}=9.25$ THz, $\omega_{K}=0.14$ THz, $\alpha=0.01$, $\beta=0.02$, $J_{x}=1$ GA///m2, $J_{y}=2$ GA///m2, and $\rho\approx 0.13$ Hz///(A///m2). The values of $\omega_{E}$, $\omega_{K}$, and $\rho$ are derived from the magnetic parameters of MnF2 [48] for a film thickness $d=3$nm and SOT efficiency $\xi=0.32$.
The phases in Eq. (10) are written as
$\phi_{1,2}^{x}$ $=$ $\tan^{-1}\left(\pm Q_{1}J_{x}\!+\!Q_{2}J_{y},\pm Q_{3}J_{x}\!+\!Q_{4}J_{y}\right),$ (13)
$\phi_{1,2}^{y}$ $=$ $\tan^{-1}\left(Q_{4}J_{x}\!\pm\!Q_{3}J_{y},-Q_{2}J_{x}\!\mp\!Q_{1}J_{y}\right),$ (14)
where $\operatorname{\tan}^{-1}(x,y)$ denotes the two-argument arctangent that returns the angle of the vector $(x,y)$, correctly resolving the quadrant [47]. In Eqs. (13) and (14), the coefficients before $J_{x,y}$, relying on $\omega$, are given by $Q_{1}=(1+\alpha\beta)\Omega_{1}\omega^{2}-[\alpha(\alpha-\beta)\omega^{2}+\Omega_{1}^{2}]\Omega_{2}$, $Q_{2}=\omega\{(1+\alpha^{2})(1+\alpha\beta)\omega^{2}+[\alpha\beta\Omega_{2}-\Omega_{1}]\Omega_{2}\}$, $Q_{3}=\omega\{(1+\alpha^{2})(\alpha-\beta)\omega^{2}+[\alpha\Omega_{1}+\beta\Omega_{2}]\Omega_{1}\}$, and $Q_{4}=(\alpha-\beta)\Omega_{2}\omega^{2}+[\alpha(1+\alpha\beta)\omega^{2}+\beta\Omega_{2}^{2}]\Omega_{1}$. Eqs. (13) and (14) show that the phase lag of the stationary forced oscillations relative to the driving SOTs depends on the amplitude ratio $J_{x}/J_{y}$ and on the driving frequency $\omega$.
The handedness of the linear modes is determined by the phase difference between the $x$- and $y$-components of $\mathbf{m}_{1,2}$. From Eqs. (13) and (14) one obtains the difference
$\Delta\phi_{1,2}\equiv\phi_{1,2}^{x}-\phi_{1,2}^{y}=\tan^{-1}\left(X_{1,2},Y_{1,2}\right),$ (15)
where $X_{1,2}=\pm(Q_{1}Q_{4}-Q_{2}Q_{3})(J_{x}^{2}-J_{y}^{2})$, and $Y_{1,2}=\pm(Q_{1}Q_{2}+Q_{3}Q_{4})(J_{x}^{2}+J_{y}^{2})+(Q_{1}^{2}+Q_{2}^{2}+Q_{3}^{2}+Q_{4}^{2})J_{x}J_{y}$.
When the SOT frequency matches the intrinsic eigenfrequencies of inertial AFMs, the amplitudes of magnetization oscillations exhibit resonant enhancement, as one can see in the resonant spectra. Fig. 2 shows the precessional and nutational spectra for different inertial relaxation times $\eta$, considering the general case $J_{x}\neq J_{y}$. The left (right) column displays spectra at negative (positive) frequencies. It is important to note that the panels (a) and (b), (c) and (d), (e) and (f), and (g) and (h) reveals an asymmetry of the spectra with respect to frequency inversion for a given $\eta$. Additionally, comparing panels (a) and (f), (b) and (e), (c) and (h), and (d) and (g), one observes that the positive-frequency spectrum of $m_{1}^{x,y}$ mirrors the negative-frequency spectrum of $m_{2}^{x,y}$, and vice versa. This can be understood from symmetry arguments: exchanging the sublattices ($\mathcal{A}_{1}^{x,y}\leftrightarrow\mathcal{A}_{2}^{x,y}$) is equivalent to time reversal ($\omega\leftrightarrow-\omega$). Therefore, the spectra are symmetric under the combined transformation $1\leftrightarrow 2$ and $\omega\leftrightarrow-\omega$.
Figure 3: (color online). (a) Resonant frequencies as a function of $\eta$. Curves are computed from Eq. (17), while markers denote peak positions extracted from the resonant spectra. A logarithmic scale is used on the frequency axis, and low-frequency regions are omitted for clarity. (b) and (c) show the dependence of peak heights on $\eta$ for precession and nutation, respectively. Curves are computed from Eqs. (20) and (19), and markers indicate values extracted from the spectra. The parameters used are the same as in Fig. 2.
Fig. 2 also shows that the precessional resonances are stronger (higher peaks) and occur at lower frequencies compared to nutational ones. However, both resonances lie in the terahertz range, due to the exchange enhancement characteristic of AFMs.
IV Frequency and magnitude of resonances
From Eqs. (11) and (12), and noting that $\alpha^{2}\ll 1$, the amplitudes of the linear modes increases at resonance, which occurs when
$\omega^{2}=\Omega_{1}\Omega_{2},$ (16)
where $\Omega_{1,2}$ are defined in Eqs. (8) and (9). Solving Eq. (16) yields the resonant frequencies, which appear in pairs ±$\pm\omega$. In the absence of spin inertia ($\eta=0$), the resonance condition gives $\omega=\pm\sqrt{\omega_{K}(2\omega_{E}+\omega_{K})}$. In the presence of inertia, the nutational frequency $\omega_{n}$ and the precessional frequency $\omega_{p}$ are given by
$\omega_{n,p}=\sqrt{\left(\frac{\kappa_{n,p}}{2\eta}\right)^{2}-\omega_{E}^{2}},$ (17)
where
$\kappa_{n,p}=\sqrt{\left(1+2\eta\omega_{E}\right)^{2}+4\eta\omega_{K}}\pm 1,$ (18)
with the upper (lower) sign corresponding to nutation (precession). The resonant frequencies exhibit strong dependence on the inertial relaxation time $\eta$, as shown in Fig. 3a.
Since the exchange interaction dominates the dynamics in AFMs, the precessional resonance is enhanced into the terahertz regime. At the same time, the nutational frequency is significantly higher than in FMs. Moreover, spin inertia induces a pronounced redshift in the resonant frequencies of both nutation and precession, as also evident in the spectra shown in Fig. 2.
In general, resonant peak heights quantify the energy stored in the forced oscillations. As illustrated in Fig. 2, these heights depend on the inertial relaxation time $\eta$ and differ for positive and negative frequencies. Therefore, it is necessary to analyze them in detail. Substituting the resonant frequencies (17) into Eqs. (11)-(12) yields explicit expressions for the peak heights, which are different for each sublattice and each component of magnetization. For the $\omega_{n}$-branch resonance, the peak heights of $x(y)$ components of sublattices $1$ and $2$ are as follows:
$\mathcal{N}_{1,2}^{x(y)}$ $=$ $\rho\frac{\eta}{\sqrt{2}\alpha\kappa_{n}}\sqrt{\frac{\kappa_{n}\!-\!2\eta\omega_{E}}{2\eta\left(\omega_{E}\!+\!\omega_{K}\right)\!+\!\kappa_{n}}}\times$ (19)
$\left|\!\sqrt{\kappa_{n}\!-\!2\eta\omega_{E}}J_{x(y)}\!\mp\!\sqrt{\kappa_{n}\!+\!2\eta\omega_{E}}J_{y(x)}\!\right|,$
For the $\omega_{p}$-branch resonance, the heights of $x(y)$ components of sublattices $1$ and $2$ are given by
$\mathcal{P}_{1,2}^{x(y)}$ $=$ $\rho\frac{\eta}{\sqrt{2}\alpha\kappa_{p}}\sqrt{\frac{\kappa_{p}\!+\!2\eta\omega_{E}}{2\eta\left(\omega_{E}\!+\!\omega_{K}\right)\!-\!\kappa_{p}}}\times$ (20)
$\left|\!\sqrt{\kappa_{p}\!-\!2\eta\omega_{E}}J_{x(y)}\!\pm\!\sqrt{\kappa_{p}\!+\!2\eta\omega_{E}}J_{y(x)}\!\right|,$
The peak heights of the negative ($-\omega_{n}$ and $-\omega_{p}$) branches are obtained by interchanging sublattices $1$ and $2$ in Eqs. (19) and (20), respectively. Throughout the derivation we assume $\alpha^{2}\!\ll\!1$, $\alpha\beta\!\ll\!1$, and $\beta^{2}\!\ll\!1$.
Eqs. (19) and (20) reveal several notable features. (i) The peak heights scale inversely with the damping parameter $\alpha$, a typical characteristic of resonant systems. (ii) The peak heights are highly sensitive to the relative magnitudes of $J_{x}$ and $J_{y}$. For the generic case $J_{x}\neq J_{y}$, as shown in Figs. 3b and 3c, the $x$- and $y$-components exhibit unequal peak heights, indicating elliptical polarization of the resonant modes. (iii) Only when $J_{x}=J_{y}$ do the $x$- and $y$-components exhibit equal peak heights, corresponding to circularly polarized modes. In this case, all four resonances at ±$\pm\omega_{n}$ and ±$\pm\omega_{p}$ exist. This contrasts with the FM case, where only two modes ($-\omega_{n}$ and $\omega_{p}$) appear, as seen in Figs. 7e and 7f. (iv) For specific ratios of $J_{x}$ and $J_{y}$, certain peaks vanish. For example, when $J_{y}/J_{x}=\sqrt{(\kappa_{n}+2\eta\omega_{E})/(\kappa_{n}-2\eta\omega_{E})}$, one finds $\mathcal{N}_{1}^{y}=0$, as indicated by the dashed curve in Fig. 3c. (v) The nutational peaks are generally much smaller than the precessional ones, as evident from comparison between Figs. 3b and 3c. In the limit $\eta=0$, one has $\kappa_{n}=2$ and thus $\mathcal{N}_{1,2}^{x}=\mathcal{N}_{1,2}^{y}=0$, indicating the absence of nutation. Meanwhile, $\kappa_{p}=0$ for $\eta=0$. Then, taking the limit $\eta\rightarrow 0$ in Eq. (20), the precessional peak heights approach
$\frac{\rho}{2\alpha}\frac{\left|\sqrt{2\omega_{E}+\omega_{K}}J_{x(y)}\pm\sqrt{\omega_{K}}J_{y(x)}\right|}{\left(\omega_{E}+\omega_{K}\right)\sqrt{\omega_{K}}},$ (21)
which corresponds to the resonant peak heights of AFM sublattices in the absence of spin inertia.
V Analyses of polarization and handedness
To characterise the resonance polarization we employ two quantities: the ellipticity, which measures the degree of polarization, and the handedness, extracted from the phase difference between $m_{x}$ and $m_{y}$. The ellipticity is defined as the aspect ratio $m_{y}/m_{x}$: it equals $1$ for circular polarization and tends to 00 or $\infty$ for linear polarization; intermediate values correspond to elliptical polarization. For the nutational modes we introduce ellipticity as $e^{n}_{1,2}=\mathcal{N}^{y}_{1,2}/\mathcal{N}^{x}_{1,2}$, and for the precessional modes $e^{p}_{1,2}=\mathcal{P}^{y}_{1,2}/\mathcal{P}^{x}_{1,2}$, where the subscripts $1,2$ label the two sublattices. Using Eqs. (19) and (20) we obtain
$e^{n}_{1,2}$ $=$ $\left|\frac{r_{n}J_{x}\mp J_{y}}{J_{x}\mp r_{n}J_{y}}\right|,$ (22)
$e^{p}_{1,2}$ $=$ $\left|\frac{r_{p}J_{x}\pm J_{y}}{J_{x}\pm r_{p}J_{y}}\right|,$ (23)
where
$r_{n,p}(\eta)=\sqrt{\frac{\sqrt{\left(1+2\eta\omega_{E}\right)^{2}+4\eta\omega_{K}}\pm 1+2\eta\omega_{E}}{\sqrt{\left(1+2\eta\omega_{E}\right)^{2}+4\eta\omega_{K}}\pm 1-2\eta\omega_{E}}}.$ (24)
In the limit $\eta\to 0$ one recovers $r_{p}=\sqrt{(2\omega_{E}+\omega_{K})/\omega_{K}}$ and $r_{n}=1$. Eqs. (22) and (23) are written for the positive-frequency branches; the corresponding expressions for negative frequencies follow by exchanging the sublattice indices $1\leftrightarrow 2$.
Figure 4: (color online). Ellipticity (upper row) and inter-component phase difference (lower row) as functions of the current ratio $J_{y}/J_{x}$ at positive frequencies for several values of $\eta$. Panels (a)-(d) present the nutational modes of sublattices 1 and 2, while (e)-(h) display the corresponding precessional modes. Solid curves are obtained from Eqs. (22), (23), and (25); symbols denote results calculated without approximations. Parameters are identical to those in Fig. 2.
The degree of polarization, characterized by the ellipticity, strongly depends on the ratio $J_{y}/J_{x}$ and the inertial relaxation time $\eta$, as shown in the upper panels of Fig. 4. Given similar behaviors for opposite signs of $J_{x,y}$ and resonant frequency, the analysis is restricted to $J_{x,y}>0$ and positive frequencies.
For the nutational mode, $\mathbf{m}_{1}$ becomes linearly polarized along the $x$-axis ($y$-axis) when $J_{y}/J_{x}=r_{n}$ ($J_{y}/J_{x}=1/r_{n}$), as indicated by the vanishing (diverging) ellipticity $e_{1}^{n}$ in Fig. 4(a). In contrast, Fig. 4(c) shows that $e_{2}^{n}$ remains finite and non-unity at these points, indicating elliptical polarization for $\mathbf{m}_{2}$. Thus, the nutational mode is linearly polarized in one sublattice and elliptically polarized in the other.
A similar feature appears in the precessional mode: at $J_{y}/J_{x}=r_{p}$ ($J_{y}/J_{x}=1/r_{p}$), $\mathbf{m}_{2}$ becomes linearly polarized along the $x$-axis ($y$-axis), while $\mathbf{m}_{1}$ remains elliptically polarized, as seen in Figs. 4(e) and 4(g).
In the limiting case where $J_{x}$ or $J_{y}$ vanishes, all modes are elliptically polarized with ellipticity equal to $r_{n,p}$ or $1/r_{n,p}$. When $J_{x}=J_{y}$, the ellipticity becomes unity, and all modes are circularly polarized. These polarization characteristics are also illustrated schematically in Fig. 5 for representative values of $J_{y}/J_{x}$, in agreement with Fig. 4.
Comparison of the ellipticity curves for different $\eta$ values reveals that increasing spin inertia enhances the deviation of $e^{n}_{1,2}$ and $e^{p}_{1,2}$ from unity, thereby increasing the ellipticity and, thus, the degree of polarization.
Refer to caption Refer to caption
Figure 5: (color online). Polarization and handedness of resonant modes for four branches (±$\pm\omega_{p}$, ±$\pm\omega_{n}$) at various values of $J_{y}/J_{x}$. The inertial relaxation time is set to $\eta=100$ fs, yielding $r_{n}=1.6855$ and $r_{p}=19.4488$ from Eq. (24). In each panel, the horizontal and vertical axes correspond to $m_{x}$ and $m_{y}$ components, respectively; axis ranges are scaled according to the larger amplitude. Curves represent the magnetization orbits, and arrows indicate the sense of rotation.
Besides ellipticity, the handedness – the sense of rotation of $\mathbf{m}_{1,2}$ around the $z$-axis – completes the description of the polarization. Let $\mathbf{m}_{\perp}=m_{k}^{x}\mathbf{e}_{x}+m_{k}^{y}\mathbf{e}_{y}$ denoting the in-plane magnetization. As outlined in Figs. 6a and 6b, a mode is righthanded if $\mathbf{e}_{z}\cdot(\mathbf{m}_{\perp}\times d\mathbf{m}_{\perp}/dt)>0$ or, equivalently, if $\omega\sin(\phi^{x}_{k}-\phi^{y}_{k})>0$ according to Eq. (10). The sign inversion corresponds to the opposite handedness.
At resonance, the phase differences defined in Eq. (15) can be simplified as
$\Delta\phi_{1,2}^{n(p)}\equiv\phi^{x}_{1,2}-\phi^{y}_{1,2}=\tan^{-1}\left(X_{1,2}^{n(p)},Y_{1,2}^{n(p)}\right),$ (25)
where, for the $\omega_{n}$($\omega_{p}$)-branch resonance,
$X_{1,2}^{n(p)}$ $=$ ±$\pm\alpha\omega_{n(p)}^{2}\left(J_{y}^{2}-J_{x}^{2}\right),$ (26)
$Y_{1,2}^{n}\!$ $=$ $\!\mp\frac{\kappa_{n}\!-\!2\eta\omega_{E}}{4\eta^{2}}\left(\sqrt{\kappa_{n}\!-\!2\eta\omega_{E}}J_{x}\!\mp\!\sqrt{\kappa_{n}\!+\!2\eta\omega_{E}}J_{y}\right)$ (27)
×$\times\left(\sqrt{\kappa_{n}\!+\!2\eta\omega_{E}}J_{x}\!\mp\!\sqrt{\kappa_{n}\!-\!2\eta\omega_{E}}J_{y}\right),$
$Y_{1,2}^{p}\!$ $=$ ±$\!\pm\frac{\kappa_{p}\!+\!2\eta\omega_{E}}{4\eta^{2}}\left(\sqrt{\kappa_{p}\!-\!2\eta\omega_{E}}J_{x}\!\pm\!\sqrt{\kappa_{p}\!+\!2\eta\omega_{E}}J_{y}\right)$ (28)
×$\times\left(\sqrt{\kappa_{p}\!+\!2\eta\omega_{E}}J_{x}\!\pm\!\sqrt{\kappa_{p}\!-\!2\eta\omega_{E}}J_{y}\right).$
where $\kappa_{n,p}$ has been defined in Eq. (18). The phase differences for the $-\omega_{n,p}$ branches follow by exchanging the sublattice indices $1\leftrightarrow 2$ in Eqs. (26)-(28). In deriving Eq. (25) we insert the resonant frequencies for $\omega$ and use the approximations $\alpha^{2}\ll 1$, $\alpha\beta\ll 1$, and $\beta^{2}\ll 1$. For positive resonant frequencies (righthanded SOTs) the mode is righthanded when $0^{\circ}<\Delta\phi_{1,2}^{n(p)}<180^{\circ}$ and lefthanded when $180^{\circ}<\Delta\phi_{1,2}^{n(p)}<360^{\circ}$.
Figure 6: (color online) (a) and (b) Definition of handedness. (c) and (d) Phase diagrams illustrating the polarization of $\mathbf{m}_{1,2}$ as functions of $J_{y}/J_{x}$ and $\eta$ for (c) nutational and (d) precessional resonances. In (c) the upper (lower) boundary is given by $J_{y}/J_{x}=r_{n}$ ($J_{y}/J_{x}=1/r_{n}$); in (d) by $J_{y}/J_{x}=r_{p}$ ($J_{y}/J_{x}=1/r_{p}$). The upper and lower circles in (d) mark $J_{y}/J_{x}=\sqrt{(2\omega_{E}+\omega_{K})/\omega_{K}}$ and $\sqrt{\omega_{K}/(2\omega_{E}+\omega_{K})}$, respectively. The tables above (c) and (d) contain the handedness in different regions for nutational and precessional resonant modes, respectively.
Analogous to the ellipticity, the phase difference $\Delta\phi_{1,2}^{n(p)}$ between the $x$- and $y$-components is governed by the ratio $J_{y}/J_{x}$ (lower panels of Fig. 4). For the nutational branch, the phase difference is $\Delta\phi_{1}^{n}=0^{\circ}$ at $J_{y}/J_{x}=r_{n}$ and $\Delta\phi_{1}^{n}=180^{\circ}$ at $J_{y}/J_{x}=1/r_{n}$ [Fig. 4(b)]. Hence $m_{1}^{x}$ and $m_{1}^{y}$ oscillate in phase at $J_{y}/J_{x}=r_{n}$ or antiphase at $J_{y}/J_{x}=1/r_{n}$, whereas $\mathbf{m}_{1}$ is linearly polarised. In the precessional branch, $\Delta\phi_{2}^{p}=180^{\circ}$ ($\Delta\phi_{2}^{p}=0^{\circ}$) at $J_{y}/J_{x}=r_{p}$ ($J_{y}/J_{x}=1/r_{p}$) [Fig. 4(h)], again indicating that $\mathbf{m}_{2}$ is linear polarized.
In the limiting case $J_{x}\to 0$ or $J_{y}\to 0$, and using $\omega_{K}\ll\omega_{E}$, in Eqs. (26)–(28), one finds $\Delta\phi\to 270^{\circ}$ for the nutational mode of $\mathbf{m}_{1}$ and the precessional mode of $\mathbf{m}_{2}$, indicating lefthanded polarization. In contrast, $\Delta\phi$ approaches $90^{\circ}$ for the nutational mode of $\mathbf{m}_{2}$ and the precessional mode of $\mathbf{m}_{1}$, corresponding to righthanded polarization. When $J_{x}=J_{y}$, all modes exhibit $\Delta\phi=90^{\circ}$, indicating right-handed circular polarization.
Using Eqs. (25)-(28), the polarization handedness of the nutational and precessional resonant modes in sublattices 1 and 2 can be determined by the sign of $\omega_{r}\sin\Delta\phi_{1,2}^{n(p)}$, where $\omega_{r}$ denotes the resonant frequency. According to this criterion, the condition $\sin\Delta\phi_{1,2}^{n(p)}=0$ defines the boundaries between regions of different handedness. These boundaries correspond precisely to $J_{y}/J_{x}=r_{n,p}$ and $J_{y}/J_{x}=1/r_{n,p}$, where $r_{n,p}$ is given in Eq. (24).
Based on this analysis, we construct the phase diagrams shown in Figs. 6(c) and 6(d), mapping the polarization types in the parameter space spanned by $\eta$ and $J_{y}/J_{x}$. The boundary curves in Fig. 6 are calculated using the approximate analytical expression (24). The shaded and unshaded regions are determined by evaluating the sign of $\omega_{r}\sin\Delta\phi_{1,2}$, where $\Delta\phi_{1,2}$ is computed from Eq. (15) using the corresponding resonant frequencies. The analytical approximations agree very well with the exact numerical results.
From the above analysis – particularly Figs. 5 and 6 – it is evident that the polarization of resonant modes can be tuned by controlling both the handedness (i.e., the sign of the frequency) and the ellipticity ($J_{y}/J_{x}$) of the alternating driving force (here, SOTs generated by alternating currents). For the sake of clarity, we restrict the discussion to positive $J_{x,y}$, such that the handedness of the driving force is solely determined by the sign of its frequency.
For both nutational and precessional modes, the polarization of all resonant modes matches that of the SOTs when the ellipticity of the drive is close to $1$ that is, when $1/r_{n}<J_{y}/J_{x}<r_{n}$ (nutation) or $1/r_{p}<J_{y}/J_{x}<r_{p}$ (precession). This corresponds to Region I in Figs. 6(c) and 6(d). In this regime, both $\mathbf{m}_{1}$ and $\mathbf{m}_{2}$ rotate righthandedly about the $+z$-axis for positive resonant frequencies, and lefthandedly for negative frequencies, consistent with the behavior shown schematically in Fig. 5.
When the ellipticity of the SOTs deviates more strongly from $1$, the polarization behavior becomes sublattice-dependent [Region II in Figs. 6(c) and 6(d)]. Specifically, for nutational modes in Region II ($J_{y}/J_{x}<1/r_{n}$ or $J_{y}/J_{x}>r_{n}$), $\mathbf{m}_{1}$ rotates lefthandedly while $\mathbf{m}_{2}$ maintains righthanded rotation at positive frequency. For precessional modes in Region II ($J_{y}/J_{x}<1/r_{p}$ or $J_{y}/J_{x}>r_{p}$), the handedness is reversed: $\mathbf{m}_{1}$ rotates righthandedly, while $\mathbf{m}_{2}$ rotates lefthandedly for positive frequency.
A comparison between the corresponding panels for positive and negative frequencies in Fig. 5 reveals that flipping the sign of the frequency is equivalent to exchanging $\mathbf{m}_{1}\leftrightarrow\mathbf{m}_{2}$ and simultaneously reversing the rotational sense. This indicates a symmetry of the resonant modes under the combined transformation: sublattice exchange $1\leftrightarrow 2$ and frequency inversion $\omega\leftrightarrow-\omega$.
In FMs, as shown in Appendix B and Fig. 8, the handedness of resonant modes is independent of the ratio $J_{y}/J_{x}$. For positive (negative) frequency, the nutational mode is lefthanded (righthanded), while the precessional mode is righthanded (lefthanded). In contrast, the situation in AFMs is more intricate. As $J_{y}/J_{x}$ crosses the critical values $r_{n(p)}$ or $1/r_{n(p)}$, the handedness of the nutational (precessional) mode is reversed for one of the sublattices. These critical points depend on the inertial relaxation time $\eta$ and are a unique feature of AFMs, absent in the FM case.
Finally, the field-like SOT [the $\beta$ term in Eq. (3)] has a negligible impact on the resonance modes. In the peak-height and phase-difference expressions, $\beta$ enters only through the factors $1+\alpha\beta$ and $1+\beta^{2}$, which satisfy $\alpha\beta\ll 1$ and $\beta^{2}\ll 1$ for typical torques induced by the spin Hall effect [42]. Numerical simulations confirm the accuracy of the resulting analytical approximation.
VI Conclusion
In this paper, we investigate resonant magnetization precession and nutation in a bilayer system composed of a uniaxial AFM and an adjacent HM layer, driven by two in-plane, mutually perpendicular alternating currents. The amplitude ratio and phase difference of these currents define the handedness and degree of polarization of the driving SOTs.
Using zero-field AFM resonant spectra and approximate analytical expressions, we show that the resonant frequencies are intrinsic, while the resonant peak heights and polarizations are sensitive to both intrinsic parameters (such as the damping factor and inertial relaxation time) and the polarization of the driving force. We further demonstrate that both nutational and precessional resonant modes always appear in pairs for arbitrary amplitude ratios of the currents – unlike in FMs, where only one nutational and one precessional mode appear when the current amplitudes are equal.
Moreover, the resonant modes in AFMs can continuously evolve through elliptical, circular, and linear polarization states by varying the amplitude ratio of the driving currents. In contrast, the resonant modes in FMs remain nearly circularly polarized. Most notably, we identify a critical current ratio at which the resonant mode becomes linearly polarized and the handedness switches across this point. This critical ratio depends on the spin inertia and the type of resonance and has no counterpart in FMs.
Our findings highlight the richer landscape of polarization and handedness in AFM resonant dynamics and suggest potential applications for AFM modes as carriers of information in spintronic devices. This work advances the fundamental understanding of inertial resonant modes and provides a basis for chirality control in spin-wave-based information processing.
VII Acknowledgments
Peng-Bin He was supported by the NSF of Changsha City (Grant No. kq2208008) and the NSF of Hunan Province (Grant No. 2023JJ30116). Ri-Xing Wang was supported by the key Program of Education Bureau of Hunan Province (Grant No.24A0494) and the Regional Joint Funds of the NSF of Hunan Province (Grant No.2024JJ7312).
Appendix A steady-state linear modes in FMs driven by SOTs
The ILLG equation of FMs driven by SOTs reads
$\frac{d\mathbf{m}}{dt}=\mathbf{m}\times\frac{d\mathcal{E}}{d\mathbf{m}}+\alpha\mathbf{m}\times\frac{d\mathbf{m}}{dt}+\eta\mathbf{m}\times\frac{d^{2}\mathbf{m}}{dt^{2}}+\boldsymbol{\tau},$ (29)
where $\mathbf{m}$ is the unit vector of magnetization. The magnetic energy arises from perpendicular magnetocrystalline anisotropy and the demagnetizing field (within the local approximation), and in frequency units it is given by
$\mathcal{E}=\omega_{K}\left(\mathbf{m}\cdot\mathbf{e}_{z}\right)^{2},$ (30)
where $\omega_{K}=\gamma_{0}H_{K}$, with $H_{K}$ being effective anisotropy field defined by $H_{K}=2K_{u}/M_{s}-\mu_{0}M_{s}$, with $K_{u}$ being the magnetocrystalline anisotropy constant, $\mu_{0}$ the vacuum susceptibility, and $M_{s}$ the saturation magnetization. The SOTs $\mathbf{\tau}$ read,
$\mathbf{\tau}=-\rho\left\{\mathbf{m}\times\left[\mathbf{m}\times\left(\mathbf{e}_{z}\times\mathbf{j}_{e}\right)\right]+\beta\mathbf{m}\times\left(\mathbf{e}_{z}\times\mathbf{j}_{e}\right)\right\},$ (31)
where $\rho$, $\beta$, and $\mathbf{j}_{e}$ are the same as those in Eq. (3).
By the same procedure as Sec. III, the steady-state linear modes are derived as
$\left(\begin{array}[]{c}m_{x}(t)\\ m_{y}(t)\end{array}\right)=\chi\left(\begin{array}[]{c}j_{x}(t)\\ j_{y}(t)\end{array}\right)$ (32)
In Eq. (32), the susceptibility reads
$\chi=\frac{\rho}{\Delta}\left(\begin{array}[]{cc}\chi_{1}&\chi_{2}\\ -\chi_{2}&\chi_{1}\end{array}\right),$ (33)
where $\chi_{1}=\Omega+i(\alpha-\beta)\omega$, $\chi_{2}=\beta\Omega+i(1+\alpha\beta)\omega$, and
$\Delta=\omega^{2}+\Omega^{2}+i\alpha\omega,$ (34)
with $\Omega=\omega_{K}-\eta\omega^{2}$. Taking the real parts of Eq. (32), the linear modes can be written explicitly as
$m_{x(y)}=\mathcal{A}_{x(y)}\cos\left(\omega t+\phi_{x(y)}\right).$ (35)
In Eq. (35), the amplitudes are expressed as
$\mathcal{A}_{x(y)}$ $=$ $\rho\sqrt{\frac{P_{1}J^{2}_{x(y)}+P_{2}J^{2}_{y(x)}+P_{3}J_{x}J_{y}}{\widetilde{\Delta}}},$ (36)
where $P_{1}=(\alpha-\beta)^{2}\omega^{2}+\Omega^{2}$, $P_{2}=(1+\alpha\beta)^{2}\omega^{2}+\beta^{2}\Omega^{2}$, $P_{3}=2(1+\beta^{2})\Omega\omega$, and
$\widetilde{\Delta}=\left[\left(\Omega+\omega\right)^{2}+\alpha^{2}\omega^{2}\right]\left[\left(\Omega-\omega\right)^{2}+\alpha^{2}\omega^{2}\right].$ (37)
The phases in Eq. (35) are given by
$\phi_{x}\!$ $=$ $\!\tan^{-1}\left(Q_{1}J_{x}\!+\!Q_{2}J_{y},Q_{3}J_{x}\!+\!Q_{4}J_{y}\right),$ (38)
$\phi_{y}\!$ $=$ $\!\tan^{-1}\left(Q_{4}J_{x}\!+\!Q_{3}J_{y},-Q_{2}J_{x}\!-\!Q_{1}J_{y}\right),$ (39)
where $\tan^{-1}(x,y)$ denotes the two-argument arctangent, which correctly identifies the quadrant of the point $(x,y)$ [47]. The frequency-dependent parameters $Q_{i}$ are defined as $Q_{1}=\Omega[(1+2\alpha\beta-\alpha^{2})\omega^{2}-\Omega^{2}]$, $Q_{2}=\omega[(1+\alpha^{2})(1+\alpha\beta)\omega^{2}-(1-\alpha\beta)\Omega^{2}]$, $Q_{3}=\omega[(1+\alpha^{2})(\alpha-\beta)\omega^{2}+(\alpha+\beta)\Omega^{2}]$, and $Q_{4}=\Omega[(2\alpha-\beta+\alpha^{2}\beta)\omega^{2}+\beta\Omega^{2}]$. From Eqs. (38) and (39), the phase difference between $m_{x}$ and $m_{y}$ is calculated as
$\Delta\phi\equiv\phi_{x}-\phi_{y}=\tan^{-1}\left(X,Y\right),$ (40)
where $X=[(\alpha-\beta)(1+\alpha\beta)\omega^{2}+\beta\Omega^{2}](J_{y}^{2}-J_{x}^{2})$, and $Y=(1+\beta^{2})\{\Omega\omega(J_{x}^{2}+J_{y}^{2})+[(1+\alpha^{2})\omega^{2}+\Omega^{2}]J_{x}J_{y}\}$.
Figure 7: (color online). Resonant spectra for various values of $\eta$. Peaks enclosed by circles (ellipses) correspond to precessional (nutational) resonances. To enhance visibility, both axes are shown on a logarithmic scale, and low-frequency regions are omitted. All spectra are calculated from Eq. (36) using the following parameters: $\omega_{K}=0.14$ THz, $\alpha=0.01$, $\beta=0.02$, $\rho\approx 0.13$ Hz///(A///m2); in (a)-(d), $J_{x}=1$ GA///m2, $J_{y}=2$ GA///m2; in (e) and (f), $J_{x}=J_{y}=1$ GA///m2. For comparison with AFMs, the values of $\omega_{K}$ and $\rho$ are chosen to match those in Fig. 2.
Appendix B resonant modes in FMs driven by SOTs
Assuming $\alpha^{2}\ll 1$, Eq. (36) implies that the resonant frequency satisfies
$\eta\omega^{2}\pm\omega-\omega_{K}=0.$ (41)
Without spin inertia ($\eta=0$), the resonant frequency is ±$\pm\omega_{K}$. In the presence of spin inertia, the resonant frequencies of nutation ±$\pm\omega_{n}$ and precession ±$\pm\omega_{p}$ are given by
$\omega_{n,p}=\frac{\sqrt{1+2\eta\omega_{K}}\pm 1}{2\eta}.$ (42)
From Eq. (36), the heights of the resonant peaks can be derived. Under the approximations $\alpha^{2}\ll 1$, $\alpha\beta\ll 1$, and $\beta^{2}\ll 1$, the peak heights are approximately equal for the $x$- and $y$-components of magnetization. Corresponding to the nutation and precession, the peak heights are written as
$\mathcal{H}_{n}^{\pm}$ $=$ $\rho\frac{\eta}{\alpha}\frac{\left|J_{x}\mp J_{y}\right|}{\sqrt{1+2\eta\omega_{K}}+1},$ (43)
$\mathcal{H}_{p}^{\pm}$ $=$ $\rho\frac{\eta}{\alpha}\frac{\left|J_{x}\pm J_{y}\right|}{\sqrt{1+2\eta\omega_{K}}-1}.$ (44)
Adopting the same approximation and substituting $\omega$ into Eq. (40), the phase differences corresponding to the ±$\pm\omega_{n}$ and ±$\pm\omega_{p}$ resonant modes are given by
$\Delta\phi_{n}^{\pm}$ $=$ $\tan^{-1}\left[\alpha\left(J_{y}^{2}-J_{x}^{2}\right),\mp\left(J_{x}\mp J_{y}\right)^{2}\right],$ (45)
$\Delta\phi_{p}^{\pm}$ $=$ $\tan^{-1}\left[\alpha\left(J_{y}^{2}-J_{x}^{2}\right),\pm\left(J_{x}\pm J_{y}\right)^{2}\right].$ (46)
Eqs. (43) and (44) imply that, for the circularly polarized SOTs ($J_{x}=J_{y}$), the nutational mode at $\omega_{n}$ and the precessional mode at $-\omega_{p}$ are nearly suppressed, as shown in Figs. 7(e) and 7(f). Meanwhile, Eqs. (45) and (46) indicate that the phase differences corresponding to the $-\omega_{n}$ and $\omega_{p}$ resonances are both $90^{\circ}$. Consequently, one finds $-\omega_{n}\sin\Delta\phi_{n}^{-}<0$ and $\omega_{p}\sin\Delta\phi_{p}^{+}>0$, indicating that only the lefthanded nutational mode and righthanded precessional mode are excited by circularly polarized SOTs. This is also illustrated in the third column of Fig. 8. These results qualitatively agree with Ref. 29, where a circularly polarized oscillating magnetic field was used to drive the system.
However, for elliptically polarized SOTs ($J_{x}\neq J_{y}$), all resonant modes – both at positive and negative frequencies – can be excited, as shown in Figs. 7(a)-(d). Moreover, all modes are nearly circularly polarized. For nutational resonances, the phase differences satisfy $180^{\circ}<\Delta\phi_{n}^{+}<360^{\circ}$ and $0^{\circ}<\Delta\phi_{n}^{-}<180^{\circ}$, leading to ±$\pm\omega_{n}\sin\Delta\phi_{n}^{\pm}<0$. Therefore, both the positive- and negative-frequency nutational modes (first and third rows of Fig. 8, respectively) are lefthanded. Similarly, for precessional resonances, one has $0^{\circ}<\Delta\phi_{p}^{+}<180^{\circ}$ and $180^{\circ}<\Delta\phi_{p}^{-}<360^{\circ}$, which results in ±$\pm\omega_{p}\sin\Delta\phi_{p}^{\pm}>0$. Consequently, both the positive- and negative-frequency precessional modes (second and fourth rows of Fig. 8) are righthanded. In summary, regardless of whether the SOT is righthanded or lefthanded, nutational resonant modes in FMs are always lefthanded, while precessional resonant modes are always righthanded.
Figure 8: (color online). Polarization and handedness of resonant modes for positive resonant frequencies and various ratios $J_{y}/J_{x}$. The inertial relaxation time is set to $\eta=100$ fs. In each panel, the horizontal and vertical axes represent the $m_{x}$ and $m_{y}$ components, respectively. The axis ranges are scaled by the larger of these two components. The curves depict the trajectories of the magnetization vector, and the arrows indicate the sense of rotation.